OpenStudy (anonymous):

Convert 5 pounds into ounces

OpenStudy (goldphenoix):

How many ounces are there in 1 pound?

OpenStudy (anonymous):

16

OpenStudy (anonymous):

\[(5\text{lbs})\cdot\frac{16\text{oz}}{1\text{lb}}=\cdots\]

OpenStudy (mathstudent55):

\( 5 ~lb \times \dfrac{16 ~oz}{1~lb} =\)

OpenStudy (goldphenoix):

Great. So if 1 pound = 16 ounces. Then 5 pounds = ? ounces?  Do you know how to cross multiply?

OpenStudy (anonymous):

I Think So

OpenStudy (goldphenoix):

So what would x be?

OpenStudy (anonymous):

80

OpenStudy (goldphenoix):

Good job. So 5 pounds is equal to 80 ounces. I hope this help! :)

OpenStudy (anonymous):

Thanxx It Did